How many grams of a 26.8% sugar solution contain 29.5 g of sugar?
1
Expert's answer
2019-09-27T04:45:38-0400

Solution.

"w = \frac{m(sugar)}{m(solution)} \times 100 \%"

"m(solution) = \frac{m(sugar)}{w} \times 100 \%"

m(solution) = 110.08 g

Answer:

m(solution) = 110.08 g